OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined. (n^3-3n)/(n^2-9)

OpenStudy (anonymous):

alright what do you know about when your equation is undefined?

OpenStudy (anonymous):

ummm...nothing?

OpenStudy (anonymous):

alright so we'll start with the basis idea that is used all the way through to calculus. is \[\frac{1}{0},\frac{25}{0},\frac{x}{0}\] computable?

OpenStudy (anonymous):

how do i tell?

OpenStudy (anonymous):

no, because nothing is divisible by 0

OpenStudy (anonymous):

well for one, when you put this into a calculator, what does it tell you?

OpenStudy (anonymous):

it says cannot divide by zero

OpenStudy (anonymous):

correct it's undefined =nothing.. so in your case you have a the same type of fraction only with variables

OpenStudy (anonymous):

so you don't have to worry about the numerator, just the denominator; you want the denominator to equal zero because in this case the problem will be undefined

OpenStudy (anonymous):

okay, n^2-9=0?

OpenStudy (anonymous):

yep do you know how to solve this equation

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

well you can either tell that it's a difference of two squares which has the equation \[(a^2-b^2)=(a+b)(a-b)\] or you can simply solve it by using algebra and adding the 9 over to right side to get \[x^2=9\] in which you take the square root of both sides

OpenStudy (anonymous):

so, x=3?

OpenStudy (anonymous):

if you did the square root way you get a positive and negative answer \[x=+3,-3\] if you did the difference method you find out that b = 3 and a = x so theequation is (x+3)(x-3)=0 which you get x+3=0 and x-3=0 x=-3,+3

OpenStudy (anonymous):

since x^2 turns any integer positive, it can be negative or positive (-3)^2=9-9=0 (3)^2=9-9=0

OpenStudy (anonymous):

I don't understand the relation between my fraction and that formula...

OpenStudy (anonymous):

the difference formula you mean?

OpenStudy (anonymous):

thoroughly confused

OpenStudy (anonymous):

I'm working on y alg 2 final...and I am so lost, confused and frustrated...

OpenStudy (anonymous):

alright you have \[n^2-9=0\] these are two perfect square. by perfect squares i mean that if you take the square root of each you get an integer so if you use \[a^2+b^2=(a+b)(a-b)\] where \[a^2=n^2,b^2=9\] you know that taking the square root you get a=n b=3 so then you just use the rest of the formula

OpenStudy (anonymous):

(n-3)(n+3)=0 now the way to make this equation zero is making one of the groups turn to zero because if you do that the whole equation will become zero... so say you make n+3=0, n would have to equal -3 right? (-3-3)(-3+3)=0 (-6)(0)=0 0=0

OpenStudy (anonymous):

does wher you put the plus or minus matter?

OpenStudy (anonymous):

for the answers? or for the equation the equation can be (a+b)(a-b) or (a-b)(a+b) if that's what you're asking.

OpenStudy (anonymous):

that's what I was asking

OpenStudy (anonymous):

yeah it'st he same as multiplying say (3) and (4) 3(4)=12 4(3)=12

OpenStudy (anonymous):

oh, okay

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No problem =]